# OODEval: Evaluating Large Language Models on Object-Oriented Design

### v1

@@ Abstract.
Recent advances in large language models (LLMs) have prompted extensive evaluations across software engineering tasks, yet existing studies largely focus on code-level performance, leaving LLMs’ capabilities in software design underexplored. To address this gap, we conduct a comprehensive empirical study evaluating 29 LLMs on object-oriented design (OOD) tasks. Given the absence of standardized benchmarks and metrics in this domain, we introduce OODEval, a manually constructed benchmark covering 50 object-oriented design tasks of varying difficulty. Moreover, we present OODEval-Human, the first human-rated OOD benchmark comprising 940 undergraduate-submitted solutions with instructor ratings. We further propose CLUE (Class Likeness Unified Evaluation), a unified metric set for assessing global and fine-grained performance in class diagram design. Using these new benchmarks and metrics, we investigate five research questions: overall correctness, comparison with humans, model dimension analysis, task feature analysis, and bad case analysis. Results show that LLMs achieve high syntactic accuracy but exhibit notable semantic deficiencies, especially in method and relationship generation. Qwen3-Coder-30B achieves the best overall performance, rivaling DeepSeek-R1 and GPT-4o, while Gemma3-4B-IT outperforms GPT-4o-Mini despite its smaller scale. Although the best LLMs nearly match the performance of average undergraduates, they remain far below the capabilities of the top humans. Parameter scale, code specialization, and instruction tuning emerge as key factors influencing LLMs’ performance, while increased design complexity and lower requirement readability degrade it. Bad case analysis reveals prevalent failure modes, such as keyword errors, missing classes or relationships, and method omissions. Overall, this study establishes a systematic evaluation framework tailored to object-oriented design tasks and provides empirical insights to guide future research, practical deployment, and educational strategies concerning the application of large language models in software design.
@@
copyright: noneccs: Software and its engineering Software design engineeringccs: General and reference Empirical studiesccs: General and reference Metricsccs: Computing methodologies Information extraction
@@ 1. Introduction
In recent years, large language models (LLMs) such as GPT(Achiam et al., 2023) and LLaMA(Touvron et al., 2023) have made remarkable advancements in the field of software engineering. To systematically assess their practical effectiveness, researchers have developed various benchmarks (Du et al., 2024; Cao et al., 2024; Ding et al., 2025; Jiang et al., 2024) for software engineering tasks and conducted empirical studies to analyze the strengths and limitations of LLMs. However, recent studies (Hu et al., 2025) indicate that existing benchmarks and empirical research are mainly focus on programming assistance and quality management tasks (77%), while requirements and design tasks represent only 8%. Moreover, the limited requirements and design tasks researches only focus primarily on isolated scenarios of requirements analysis and verification, which lacks systematic evaluation of LLMs’ capabilities from requirements to design, thus leaving it unclear how LLMs perform in software design scenarios.
@@
  • CLUE Metrics: A novel OOD correctness evaluation metric, containing both an overall performance metric and four local performance metrics. These metrics are confirmed to be high correlation with human ratings on OODEval-Human.
@@ 2.1. Existing Datasets for OOD Generation
OODEval Model-level 50 8 $\pm$ 3 19 $\pm$ 10 17 $\pm$ 16 9 $\pm$ 4 237 $\pm$ 101 JSON/PNG
@@ 2.3. Existing Empirical Studies for OOD Generation
In recent years, researchers have actively explored the potential of Large Language Models (LLMs) in automated domain and conceptual modeling, with growing attention on their ability to generate UML class diagrams. Representative works include (Chen et al., 2023b), which manually compared GPT-3.5 and GPT-4 on 10 cases across classes, attributes, and relationships; (Cámara et al., 2023), which examined ChatGPT’s performance on UML diagrams and OCL constraints using 8 cases; (De Bari et al., 2024), which proposed a multidimensional (syntax, semantics, practicality, correctness) evaluation framework comparing GPT-4 with human designs; and (Calamo et al., 2025), which introduced automated evaluation and analyzed the impact of model scale and task complexity. These studies have deepened understanding of LLMs in object-oriented design and provided valuable evaluation paradigms. However, current research has notable limitations: (1) small-scale experiments (typically $\leq 10$ cases) and narrow model coverage restrict generalizability; (2) reliance on single-expert designs or unvalidated automated metrics introduces subjectivity and uncertainty; (3) insufficient analysis of defect types/frequencies and near-absence of class method evaluation.
@@ 3.1. OODEval: Benchmark for OOD Evaluation
3.1.1. Construction Procedure.
The OODEval benchmark is designed to evaluate the ability of LLMs to generate class diagrams design from natural language requirements. The primary challenge in constructing a high-quality OOD dataset lies in the collection of software requirements. Publicly available software requirements are already scarce, and resources that provide both requirements and the corresponding design class diagrams are even more limited. Existing datasets (e.g., OOPEval(Wang et al., 2024) and RepoClassBench(Deshpande et al., 2024)) automatically generate requirement descriptions from code. However, the absence of a rigorous requirement modeling process often leads to incomplete or inaccurate results, thereby reducing the quality and applicability of these datasets. To address these limitations, we adopted a multi-source collection strategy. First, we reviewed literature published over the past decade on mapping from requirements to design and retrieved 20 usable open source OOD datasets from appendices and online repositories. Second, leveraging GitHub’s native support for the Mermaid modeling language, we performed a keyword search using the query “mermaid classdiagram language:Markdown AND requirement path:*.md” and obtained 13 valid instances that contained both requirements and class diagrams. Third, we curated 17 pairs of requirements and designs from the project archives of undergraduate OOD courses. Through multiple rounds of collection and screening, we obtained 50 instances, with task IDs encoded from 1 to 50, each containing high-quality requirements and design specifications.
@@
Figure 1. An Example of Input and Output Format in OODEval Benchmark.
@@ 3.2. OODEval-Human: Undergraduate OOD Solutions with Instructor Ratings
3.2.3. Key Characteristics.
Table 4 presents the statistics of the OODEval-Human dataset, which consists of four undergraduate OOD tasks, each associated with OODEval by its Task ID number. Each task includes over 200 student submissions, totaling 940 submissions, with individual scores ranging from 32 to 100. The average scores for the tasks range from 79.3 to 82.8, with an overall average of 80.62. This provides a sufficient sample size and score distribution for evaluating the performance of the CLUE metric. The difficulty levels of these four tasks in OODEval are moderate to hard, making them suitable for comparing the OOD capabilities of LLMs with human performance on more complex tasks.
@@ 4. CLUE: Class Likeness Unified Evaluation Metric
This section introduces a novel evaluation metric to assess the correctness of class diagram designs. After presenting preliminary definitions, we introduce the CLUE metrics, followed by hyperparameter optimization to determine optimal weights. The metrics’ efficacy and superiority are then demonstrated by comparing CLUE scores with human instructor ratings on the OODEval-Human dataset, achieving significantly higher correlation coefficients than existing baselines.
@@
The CLUE metric is designed to assess the similarity between a candidate class diagram and a reference class diagram. A pivotal distinction from other methods lies in the necessity of a designated reference model, rather than evaluating any arbitrary pair of models; this difference will be elaborated upon subsequently. We commence by defining the Class Diagram Model, as presented in Definition 4.1.
@@
A class diagram model is formally defined as a four-tuple $M=(C,A,M,R)$, where:
  • $C=\{c_{1},c_{2},\dots,c_{n}\}$ is a finite set of $n$ classes.
  • For each class $c\in C$, $A(c)=\{a_{1}(c),a_{2}(c),\dots,a_{p}(c)\}$ denotes its set of attributes, where each attribute is a pair $a=(\textit{name},\textit{type})$.
@@
  • Similarly, $M(c)=\{m_{1}(c),m_{2}(c),\dots,m_{q}(c)\}$ denotes its set of methods, where each method is a triple $m=(\textit{name},\textit{returnType},P)$, and $P=\{p_{1},\dots,p_{r}\}$ is the set of parameters, with each parameter being a pair $p=(\textit{name},\textit{type})$.
@@
  • $R=\{r_{1},r_{2},\dots,r_{s}\}$ is the set of relationships, where each relationship is a four-tuple $r=(\textit{r\_type},\textit{c\_begin},$ $\textit{c\_end},\textit{label})$. Here, $\textit{r\_type}\in\mathbb{T}=\{\text{AS},\text{AG},\text{CO},\text{DE},\text{GE},\text{RE}\}$ denotes the relationship type, with AS = association, AG = aggregation, CO = composition, DE = dependency, GE = generalization, and RE = realization; c_begin and c_end represent the source and target classes, respectively, thereby determining the direction; and $\textit{label}=(\textit{from},\textit{to})$ defines the multiplicity constraints at each end.
Computing the similarity between two models presents significant challenges, particularly in matching their elements when the names differ despite semantic equivalence. For instance, the classes ”Teacher” and ”Tutor” may represent substantially the same entity but bear distinct names—a common and permissible occurrence in the design domain. Consequently, the primary issue to address is identifying the optimal matching between elements of the candidate model and the reference model. In CLUE, we formulate this as an optimization problem, as defined in Definition 4.2.
@@
Let $ES\in[0,1]^{n\times m}$ be a similarity matrix of size $n\times m$, where each entry $ES[i][j]\in[0,1]$ represents the similarity between the $i$-th element of a reference set and the $j$-th element of a candidate set. The optimal matching similarity $Sim(ES)\in[0,1]$ is defined as:
(1) $Sim(ES)=\begin{cases}0.0,&\text{if }m=0\\ 1.0,&\text{if }n=0\\ \frac{1}{n}\max_{\pi:[1..n]\to[1..m]}\sum_{i=1}^{n}ES[i][\pi(i)],&\text{if }m\geq n>0\\ \frac{1}{n}\max_{\pi^{\prime}:[1..m]\to[1..n]}\sum_{j=1}^{m}ES[\pi^{\prime}(j)][j],&\text{if }n>m>0\end{cases}$
where, $\pi$ and $\pi^{\prime}$ represent injective mappings, which map from the smaller set to the larger set.
@@
Prior to implementing the optimal matching, a similarity matrix $ES$ must be constructed. Each entry in the $ES$ matrix is derived by computing the semantic similarity between two strings, as defined in Definition 4.3. Specifically, we employ the CodeBERT (Feng et al., 2020) model to generate embedding representations for $s_{1}$ and $s_{2}$, and subsequently calculate the string semantic similarity using the cosine similarity function. Furthermore, we examine the impact of alternative underlying embedding models on CLUE’s performance in Section 4.5.
Definition 4.0 (String Semantic Similarity).
Let $\sigma(s_{1},s_{2})\in[0,1]$ be a similarity function to calculate the semantic similarity between the string $s_{1}$ and the string $s_{2}$.
In class relationship designing, multiplicity constraints between two classes are typically denoted by numeric or symbolic labels at each end of the relationship. However, the string semantic similarity measure introduced earlier (Definition 4.3) fails to adequately capture the similarity of these constraints. To address this limitation, we define relationship multiplicity similarity (Definition 4.4) for any pair of class relationships as follows.
@@
Given a reference model $M_{1}$ and a candidate model $M_{2}$, for a relationship pair $(r_{1},r_{2})\in R_{1}\times R_{2}$, the relationship multiplicity similarity is:
(2) $\text{sim}_{rq}(r_{1},r_{2})=\begin{cases}0.5\cdot\text{match}(r_{1}.\text{label.from},r_{2}.\text{label.from})\\ +0.5\cdot\text{match}(r_{1}.\text{label.to},r_{2}.\text{label.to}),&\text{if }r_{1}.\text{r\_type},r_{2}.\text{r\_type}\in\{\text{AS},\text{AG},\text{CO}\}\\ 1.0,&\text{if }r_{1}.\text{r\_type},r_{2}.\text{r\_type}\in\{\text{DE},\text{GE},\text{RE}\}\\ 0.0,&\text{otherwise}\end{cases}$
where $\text{match}(q_{1},q_{2})=1$ if $q_{1}=q_{2}$ or both $q_{1}$ and $q_{2}$ contain “*”, “many”, “much”, or “multi” (matched via regular expression), and 0 otherwise.
Definition 4.5 defines the similarity between the attribute set and method set of a class pair from the reference and candidate models, computed via optimal matching to maximize the similarity score. Similarly, Definition 4.6 defines the parameter set similarity of a method pair from the reference and candidate model, also obtained through optimal matching to yield the maximum similarity score.
Definition 4.0 (Attribute/Method Set Similarity).
Given a reference model $M_{1}$ and a candidate model $M_{2}$, for a class pair $(c_{1},c_{2})$ where $c_{1}$ has $p_{1}$ attributes and $c_{2}$ has $p_{2}$ attributes, the attribute set similarity between $c_{1}$ and $c_{2}$ is defined as $\text{sim}_{ca}(c_{1},c_{2})=Sim(ES_{ca})$, where $ES_{ca}\in\mathbb{R}^{p_{1}\times p_{2}}$ is constructed as follow:
(3) $ES_{ca}[k][l]=w_{at}\cdot\sigma(a_{1}^{(k)}.\text{type},a_{2}^{(l)}.\text{type})+w_{an}\cdot\sigma(a_{1}^{(k)}.\text{name},a_{2}^{(l)}.\text{name})$
Similarly, if $c_{1}$ has $q_{1}$ methods and $c_{2}$ has $q_{2}$ methods, the method set similarity between $c_{1}$ and $c_{2}$ is defined as $\text{sim}_{cm}(c_{1},c_{2})=Sim(ES_{cm})$, where $ES_{cm}\in\mathbb{R}^{q_{1}\times q_{2}}$ is constructed as follows:
$ES_{cm}[k][l]$ $=w_{mt}\cdot\sigma(m_{1}^{(k)}.\text{returnType},m_{2}^{(l)}.\text{returnType})+w_{mn}\cdot\sigma(m_{1}^{(k)}.\text{name},m_{2}^{(l)}.\text{name})$
(4) $\quad+w_{mp}\cdot\text{sim}_{mp}(m_{1},m_{2})$
Definition 4.0 (Parameter Set Similarity).
Given a reference model $M_{1}$ and a candidate model $M_{2}$, for a method pair $(m_{1},m_{2})$ where $m_{1}$ has $r_{1}$ parameters and $m_{2}$ has $r_{2}$ parameters, the method parameter similarity $\text{sim}_{mp}(m_{1},m_{2})$ is defined as $Sim(ES_{mp})$, where the matrix $ES_{mp}\in\mathbb{R}^{r_{1}\times r_{2}}$ defined as:
(5) $ES_{mp}[k][l]=w_{pt}\cdot\sigma(p_{1}^{(k)}.\text{type},p_{2}^{(l)}.\text{type})+w_{pn}\cdot\sigma(p_{1}^{(k)}.\text{name},p_{2}^{(l)}.\text{name})$

### v2

@@ Abstract.
Recent advances in large language models (LLMs) have prompted extensive evaluations throughout software engineering tasks, yet existing studies primarily focus on code-level performance, leaving LLMs’ capabilities in software design underexplored. This gap stems mainly from the absence of standardized benchmarks and automatic evaluation metrics in this domain. To fill these gaps, this work introduces a novel evaluation framework for object-oriented design (OOD) tasks, comprising new benchmarks and metrics. Specifically, we first present OODEval, a manually constructed benchmark covering 50 OOD tasks categorized into three difficulty levels. To facilitate comparisons with human performance and the evaluation of our new metrics, we introduce OODEval-Human, the first human-rated OOD benchmark with 940 undergraduate-submitted solutions with instructor ratings. We further propose CLUE (Class Likeness Unified Evaluation), an automatic metric set for assessing global and fine-grained OOD performance. Leveraging this framework, we conduct a comprehensive empirical study evaluating 29 LLMs on OOD tasks, investigating five research questions: overall correctness, comparison with humans, model dimension analysis, task feature analysis, and bad case analysis. Results reveal that LLMs attain high syntactic accuracy but suffer from semantic deficiencies, particularly in method and relationship generation. Top-performing LLMs approach average undergraduate levels but lag behind expert humans. Key influencers include parameter scale, code specialization, and instruction tuning, while design complexity and requirement readability hinder performance. Common failure modes involve omissions and inaccuracies in classes, methods, and relationships. Overall, this study establishes a systematic evaluation framework for OOD tasks and offers empirical insights to inform future research, deployment, and education on LLMs in software design.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2026; Woodstock, NYisbn: 978-1-4503-XXXX-X/2026/06ccs: Software and its engineering Software design engineeringccs: Software and its engineering Design languagesccs: General and reference Empirical studiesccs: General and reference Metricsccs: Computing methodologies Information extraction
@@ 1. Introduction
In recent years, large language models (LLMs) such as GPT(Achiam et al., 2023) and LLaMA(Touvron et al., 2023) have made remarkable advancements in the field of software engineering. To systematically assess their practical effectiveness, researchers have developed various benchmarks (Du et al., 2024; Cao et al., 2024; Ding et al., 2025; Jiang et al., 2024) for software engineering tasks and conducted empirical studies to analyze the strengths and limitations of LLMs. However, recent studies (Hu et al., 2026) indicate that existing benchmarks and empirical research are mainly focus on programming assistance and quality management tasks (77%), while requirements and design tasks represent only 8%. Moreover, the limited requirements and design tasks researches only focus primarily on isolated scenarios of requirements analysis and verification, which lacks systematic evaluation of LLMs’ capabilities from requirements to design, thus leaving it unclear how LLMs perform in software design scenarios.
@@
In summary, this paper makes the following contributions:
@@
  • CLUE Metrics: A novel OOD automatic evaluation metric set for assessing global and fine-grained OOD performance. These metrics are confirmed to be high correlation with human ratings on OODEval-Human.
@@ 2.1. Existing Datasets for OOD Generation
OODEval 2026 Model-level 50 8 $\pm$ 3 19 $\pm$ 10 17 $\pm$ 16 9 $\pm$ 4 237 $\pm$ 101 JSON/PNG
@@ 2.3. Existing Empirical Studies for OOD Generation
In recent years, researchers have actively explored the potential of LLMs in automated modeling, such as generating UML class diagrams. Chen et al. (Chen et al., 2023b) manually compared GPT-3.5 and GPT-4 on 10 cases across classes, attributes, and relationships; Cámara et al. (Cámara et al., 2023) examined ChatGPT’s performance on UML diagrams and OCL constraints using 8 cases; De Bari et al. (De Bari et al., 2024) proposed a multidimensional (syntax, semantics, practicality, correctness) evaluation framework comparing GPT-4 with human designs; and Calamo et al. (Calamo et al., 2025), which introduced automated evaluation and analyzed the impact of model scale and task complexity. These studies have deepened understanding of LLMs in object-oriented design and provided valuable evaluation paradigms. However, current research has notable limitations: (1) small-scale experiments (typically $\leq 10$ cases) and narrow model coverage restrict generalizability; (2) reliance on single-expert designs or unvalidated automated metrics introduces subjectivity and uncertainty; (3) insufficient analysis of defect types/frequencies and the near-absence of class method evaluation.
@@ 3.1. OODEval: Benchmark for OOD Evaluation
The OODEval benchmark is designed to evaluate the ability of LLMs to generate class diagrams design from natural language requirements. The primary challenge in constructing a high-quality OOD dataset lies in the collection of software requirements. Publicly available software requirements are already scarce, and resources that provide both requirements and the corresponding design class diagrams are even more limited. Existing datasets (e.g., OOPEval(Wang et al., 2024) and RepoClassBench(Deshpande et al., 2024)) automatically generate requirement descriptions from code. However, the absence of a rigorous requirement modeling process often leads to incomplete or inaccurate results, thereby reducing the quality and applicability of these datasets. To address these limitations, we adopted a manual collection strategy drawing from diverse sources: (1) leveraging GitHub’s native support for the Mermaid modeling language, we performed a keyword search using the query “mermaid classdiagram language:Markdown AND requirement path:*.md” to identify Markdown files containing keywords of ”mermaid”, ”classdiagram”, and ”requirement”; (2) surveying literature on requirements-to-design mappings and extracting usable open-source OOD datasets from appendices and repositories; and (3) compiling requirements and designs from archived projects in university OOD courses. To avoid data leakage, we restricted collection to data created after 2025, as the latest training data cutoff date of the LLMs studied in this work is 2024 (see Table 9). Furthermore, we retained only those instances that contain complete requirements together with their corresponding class diagrams. This process ultimately yielded 50 high-quality instances (13 from GitHub, 20 from academic literature, and 17 from university courses), each accompanied by high-quality requirements and design specifications.
@@
Figure 1. An Example of Input and Output Format in OODEval Benchmark.
@@ 3.2. OODEval-Human: Undergraduate OOD Solutions with Instructor Ratings
3.2.3. Key Characteristics.
Table 4 presents the statistics of the OODEval-Human dataset, which consists of four undergraduate OOD tasks, each associated with OODEval by its Task ID number. Each task includes over 200 student submissions, totaling 940 submissions, with individual scores ranging from 32 to 100. The average scores for the tasks range from 79.3 to 82.8, with an overall average of 80.62. The CLUE metric serves as an indicator for quantifying the similarity between two designs; thus, it is independent of requirement diversity. Consequently, the distinct solutions submitted by each student ensure a sufficient sample size and diverse score distribution, thereby enabling robust evaluation of the CLUE metric’s performance, even though OODEval-Human comprises four tasks. Furthermore, the difficulty levels of these four tasks in OODEval are moderate to hard, making them well-suited for benchmarking the OOD capabilities of LLMs against human performance on more challenging tasks.
@@ 4. CLUE: Class Likeness Unified Evaluation Metric
This section introduces a novel evaluation metric for assessing the correctness of class diagram designs. First, we present the preliminary definitions required for the formulation of CLUE. We then describe the computation of the CLUE metric, followed by a hyperparameter optimization procedure to determine the optimal weights. To evaluate the effectiveness of CLUE, we adopt commonly used metric evaluation practices reported in prior work (Evtikhiev et al., 2023; Zhou et al., 2023; Xu et al., 2024), which assess the correlation between automatic metrics and human preferences; a higher correlation indicates a better metric. In this study, we compare CLUE scores with instructor ratings on the OODEval-Human dataset and observe significantly higher correlation coefficients than those achieved by other metrics.
@@
The CLUE metric is specifically devised to measure the degree of structural and semantic similarity between a candidate class diagram and a corresponding reference class diagram. Accordingly, we begin by formally defining the concept of a Class Diagram Model, as articulated in Definition 4.1.
@@
A class diagram model is formally defined as a tuple $\mathcal{D}=(C,R),$ where:
@@
  • $C=\{c_{1},c_{2},\dots,c_{n}\}$ is a finite set of classes, and each class $c=(name,A,M),$ consists of a class name,a finite set of attributes $A$, and a finite set of methods $M$. An attribute is a pair $a=(\textit{name},\textit{type})$, where both elements are strings. A method is a triple $m=(\textit{name},\textit{returnType},P)$, where name and returnType are strings, and $P=\{p_{1},\dots,p_{k}\}$ is a finite set of parameters, each parameter being a pair $p=(\textit{name},\textit{type})$ of strings.
@@
  • $R=\{r_{1},r_{2},\dots,r_{s}\}$ is a finite set of relationships, where each relationship is defined as $r=(\textit{type},\textit{begin},\textit{end},\textit{label})$. Here, $\textit{type}\in\mathbb{T}=\{\text{AS},\text{AG},\text{CO},\text{DE},\text{GE},\text{RE}\}$ denotes the relationship type (association, aggregation, composition, dependency, generalization, and realization). $\textit{begin},\textit{end}\in\{1,\dots,|C|\}$ denote the indices of the source and target classes in $C$, respectively. And $\textit{label}=(\textit{from},\textit{to})$ specifies the multiplicity constraints at both ends.
Computing the similarity between two Class Diagram Models poses substantial challenges, primarily arising from the fact that semantically equivalent model elements frequently exhibit differing lexical names. For example, the classes Teacher and Tutor may denote essentially the same conceptual entity despite their distinct labels, which is a common phenomenon in software design practice. When the elements under comparison consist solely of strings (e.g., class names, attribute names, or operation signatures), semantic similarity can be effectively captured by computing the cosine similarity between their corresponding embedding vectors, as formalized in Definition 4.2. In this work, we adopt the CodeBERT model (Feng et al., 2020) as the embedding function to generate high-quality contextual embeddings for input strings $s_{1}$ and $s_{2}$. This choice is motivated by CodeBERT’s superior performance in capturing underlying semantics, as demonstrated through an ablation study evaluating various underlying semantic models (see Section 4.5).
Definition 4.0 (String Semantic Similarity).
Let $\mathbf{e}:\mathcal{S}\to\mathbb{R}^{d}$ be an embedding function that maps strings to $d$-dimensional real vectors, where $\mathcal{S}$ denotes the space of strings. The semantic similarity $\sigma(s_{1},s_{2})\in[0,1]$ between strings $s_{1},s_{2}\in\mathcal{S}$ is given by normalized cosine similarity between their embedding vectors:
(1) $\sigma(s_{1},s_{2})=0.5\cdot(\frac{\mathbf{e}(s_{1})^{\top}\mathbf{e}(s_{2})}{\|\mathbf{e}(s_{1})\|_{2}\|\mathbf{e}(s_{2})\|_{2}}+1)$
where $\cdot^{\top}$ denotes the transpose, and $\|\cdot\|_{2}$ is the Euclidean norm.
When comparing two sets of model elements (e.g., the attribute sets of a reference class and a candidate class), the task becomes significantly more challenging. Simple exact name matching is clearly insufficient, as semantically equivalent attributes may carry different names. To address this, we first compute the pairwise similarities between every element of the reference set and every element of the candidate set, thereby constructing the Element Set Similarity Matrix as formally defined in Definition 4.3. Subsequently, we identify an optimal matching $\pi$ between the two sets that maximizes the sum of the similarities of the matched pairs (see Definition 4.4).
Definition 4.0 (Element Set Similarity Matrix).
Let $A=\{a_{1},\dots,a_{n}\}$ and $B=\{b_{1},\dots,b_{m}\}$ be two element sets. Each element $a_{i}\in A$ and $b_{j}\in B$ is represented by a $d$-tuple $(c_{1},\dots,c_{d}).$ Let $\boldsymbol{\mathrm{sim}}=(\mathrm{sim}_{1},\dots,\mathrm{sim}_{d})$ denote a collection of similarity functions, and let $\mathbf{w}=(w_{1},\dots,w_{d})$ be a weight vector satisfying $w_{k}\geq 0$ and $\sum_{k=1}^{d}w_{k}=1$. The element set similarity matrix between $A$ and $B$ is defined as $ES\in[0,1]^{n\times m}$, where each entry $ES(i,j)$ is computed as:
(2) $ES(i,j)=\sum_{k=1}^{d}w_{k}\cdot\mathrm{sim}_{k}\bigl(c_{k}(a_{i}),c_{k}(b_{j})\bigr),\quad i=1,\dots,n,\;j=1,\dots,m.$
where each similarity function $\mathrm{sim}_{k}(\cdot,\cdot)$ maps two strings (Definition 4.2), two elements (Definition 4.5), or two element sets (Definition 4.4) to a value in $[0,1]$, depending on the structural type of the $k$-th component.
@@
Let $R=\{r_{1},r_{2},\dots,r_{n}\}$ be the reference set and $C=\{c_{1},c_{2},\dots,c_{m}\}$ be the candidate set, where $n>0$ and $m>0$. Let $ES\in[0,1]^{n\times m}$ be the similarity matrix between $R$ and $C$. Each entry $ES(i,j)\in[0,1]$ represents the similarity between the $i$-th element of $R$ and the $j$-th element of CC(see Definition 4.3). The optimal matching similarity $OptMatch(R,C)\in[0,1]$ is then defined as:
(3) $OptMatch(R,C)=\begin{cases}\frac{1}{n}\max\limits_{\pi}\left\{\sum_{i=1}^{n}ES(i,\pi(i))\right\},&\text{if }m\geq n>0\\ \frac{1}{n}\max\limits_{\pi^{\prime}}\left\{\sum_{j=1}^{m}ES(\pi^{\prime}(j),j)\right\},&\text{if }n>m>0\end{cases}$
where $\pi(\textperiodcentered)$ is an injective function from $R$ to $C$, and $\pi^{\prime}(\textperiodcentered)$ is an injective function from $C$ to $R$.
@@
In class relationship designing, multiplicity constraints between two classes are typically denoted by numeric or symbolic labels at each end of the relationship. However, the string semantic similarity measure introduced earlier in Definition 4.2 fails to adequately capture the similarity of these constraints. To address this limitation, we define relationship multiplicity similarity for any pair of class relationships, as a domain-informed heuristic grounded in UML semantics, as shown in Definition 4.5.
@@
Let $\mathbb{T}_{1}=\{\text{AS},\text{AG},\text{CO}\}$ and $\mathbb{T}_{2}=\{\text{DE},\text{GE},\text{RE}\}$, where $\mathbb{T}_{1}\cup\mathbb{T}_{2}=\mathbb{T}$. Given two relationship multiplicity labels $\ell_{1}=(from_{1},to_{1})$ and $\ell_{2}=(from_{2},to_{2})$, and their corresponding relationship types $t_{1}$ and $t_{2}$, the relationship multiplicity similarity is defined as:
(4) $\operatorname{sim}_{rm}(\ell_{1},\ell_{2}\mid t_{1},t_{2})=\begin{cases}0.5\cdot\bigl(\operatorname{match}(from_{1},from_{2})+\operatorname{match}(to_{1},to_{2})\bigr),&\text{if }t_{1},t_{2}\in\mathbb{T}_{1},\\[2.84526pt] 1.0,&\text{if }t_{1},t_{2}\in\mathbb{T}_{2},\\[2.84526pt] 0.0,&\text{otherwise}.\end{cases}$
where $\operatorname{match}(q_{1},q_{2})=1$ if $q_{1}=q_{2}$ or both $q_{1}$ and $q_{2}$ match “*”, “many”, “much”, or “multi” via regular expression, and 0 otherwise.